You’re trying to choose between two different investments, both of which have up-front costs of $102,000. Investment G returns $167,000 in 7 years. Investment H returns $287,000 in 14 years. Calculate the rate of return for each these investments. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)Youre trying to choose between two different investments, both of which have up-front costs of $102,000. Investment G return
Answer #1


solution: Investment G FV=PV*(1+i)^n 167000=102000*(1+i)^7 (1+i)^7=167000/102000 (1+i)^7=1.637255 1+i=1.637255^(1/7) 1+i=1.07Investment H FV=PV*(1+i)^n 287000=102000*(1 +i)^14 (1+i)^14=287000/102000 (1+i)^14=2.813725 1+1=2.813725^(1/14) 1+1=1.076692